A boy has a lemonade stand with a cash box that has 28 coins in it. If the coins are made up of nickles and half dollars and have a value of $7.25, how many of each coins does he have in the cash box?
Answer by algebrahouse.com(1659)

x = number of nickels
28 - x = number of half-dollars

0.05x + 0.5(28 - x) = 7.25 {value of coin times number of coins equals total value}
0.05x + 14 - 0.5x = 7.25 {used distributive property}
-0.45x + 14 = 7.25 {combined like terms}
-0.45x = -6.75 {subtracted 14 from each side}
x = 15 {divided each side by -0.45}
28 - x = 13 {substituted 15, in for x, into 28 - x}

15 nickels
13 half-dollars

To ask a question, visit: www.algebrahouse.com